OpenStudy (anonymous):

@horsegirl27 @Nnesha @jagr2713 @Godlovesme @CausticSyndicalist

OpenStudy (godlovesme):

hey @LuLu152 what's ur question?

OpenStudy (anonymous):

whats your question

OpenStudy (anonymous):

here

jagr2713 (jagr2713):

all i did was add 9+5=14 and 5+2=7 so i got C

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

not C

OpenStudy (anonymous):

@Nnesha do you know

Nnesha (nnesha):

\[ \frac{ 1 }{ 5 }x+5 + \frac{ 1 }{ 9 }x + 2\]

Nnesha (nnesha):

5+2 is ??

OpenStudy (anonymous):

7

Nnesha (nnesha):

so yes and what is \[\frac{ 1}{ 5 }x + \frac{ 1}{9 }x\] can you combine this ??

OpenStudy (anonymous):

umm idk sorry @Nnesha

Nnesha (nnesha):

you can combine them bec they both have same variable and same exponent

OpenStudy (anonymous):

ok

Nnesha (nnesha):

okay so you just have to write that together they don't want u to add it so what is your answer suppose to bw

OpenStudy (anonymous):

idk im not smart

OpenStudy (anonymous):

sorry

Nnesha (nnesha):

yes you are !!! u have to learn iit to become super smart you don't this because you are learning it yet

Nnesha (nnesha):

so answer is just there they don't want u to add 1/5x and 1/9 x just write them together and add 2 and 5

OpenStudy (anonymous):

r are right but i need to know the answers faster dont give direct answers but can u go a little faster

OpenStudy (anonymous):

its not C

Nnesha (nnesha):

prettty much i told u everything

OpenStudy (anonymous):

but is it A, B, or C

Nnesha (nnesha):

WRITE 1/5xand 1/9x together or add 5+2

Nnesha (nnesha):

5 and 2 both doesn't have any variable so you can combine them what is 5+2 ?? is it 3??

OpenStudy (anonymous):

B

Nnesha (nnesha):

yes right